OpenStudy (anonymous):

The square has a perimeter of 42, solve for x, if the perimeter were greater than 42 what would we know about the value of x

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

 Now if perimeter was greater than 42 then? Even the side qould be greater than 10.5 Understood? :)

OpenStudy (anonymous):

On the diagram it has x + 3 on one side what does that mean

OpenStudy (akashdeepdeb):

That means the sides value is x+3 Ignore my diagram then... Perimeter = 4*(x+3) = 42 4x + 12 = 42 4x = 30 x = 7.5 Understood? :)

OpenStudy (anonymous):

ok thanks

OpenStudy (akashdeepdeb):

:)